A person wishes to mix coffee worth $12 per lb with coffee worth $6 per lb to get 90 lb of a mixture worth $8 per lb. How many pounds of the $12 and the $6 coffees will be needed?
Answer by fractalier(6550)

Let x be the amount used of the $12 coffee.
Then 90-x is the amount used of the $6 coffee.
The set up looks like this
12x + 6(90-x) = 8(90)
which we finish as
12x + 540 - 6x = 720
6x + 540 = 720
6x = 180
x = 30 lbs of $12
90-x = 60 lbs of $6